# Are LLMs effective psychological assessors? Leveraging adaptive RAG for interpretable mental health screening through psychometric practice

### v1

@@ Abstract
In psychological practice, standardized questionnaires serve as essential tools for assessing mental constructs (e.g., attitudes, traits, and emotions) through structured questions (aka items). With the increasing prevalence of social media platforms where users share personal experiences and emotions, researchers are exploring computational methods to leverage this data for rapid mental health screening. In this study, we propose a novel adaptive Retrieval-Augmented Generation (RAG) approach that completes psychological questionnaires by analyzing social media posts. Our method retrieves the most relevant user posts for each question in a psychological survey and uses Large Language Models (LLMs) to predict questionnaire scores in a zero-shot setting. Our findings are twofold. First we demonstrate that this approach can effectively predict users’ responses to psychological questionnaires, such as the Beck Depression Inventory II (BDI-II), achieving performance comparable to or surpassing state-of-the-art models on Reddit-based benchmark datasets without relying on training data. Second, we show how this methodology can be generalized as a scalable screening tool, as the final assessment is systematically derived by completing standardized questionnaires and tracking how individual item responses contribute to the diagnosis, aligning with established psychometric practices 111 Code available at the following link: github.com/Fede-stack/Adaptive-RAG-for-Psychological-Assessment.
@@
Federico Ravenda1,2, Seyed Ali Bahrainian3,4, Andrea Raballo1, Antonietta Mira1,5, Noriko Kando2 federico.ravenda@usi.ch, bahrainian@brown.edu andrea.raballo@usi.ch, antonietta.mira@usi.ch, kando@nii.ac.jp 1Università della Svizzera italiana, 2National Institute of Informatics 3Brown University, 4University of Tübingen, 5Insubria University
@@ 1 Introduction
According to the World Health Organization (WHO), one in seven adolescents experiences a mental health disorder Wiederhold (2022), with depression, anxiety, and behavioral disorders leading among young people. Following COVID-19, mental health conditions surged, with depressive disorders increasing by 28% in 2020 Kieling et al. (2011); Winkler et al. (2020). Given the extent of this need, the WHO Special Initiative for Mental Health prioritizes improving and expanding access to quality mental health interventions and services as a key strategic goal.222 https://www.who.int/publications/i/item/9789240049338
@@
In this work, we focus on two widely used standardized psychological questionnaires, specifically the BDI-II Beck et al. (1996) for depression screening and the Self-Harm Inventory (SHI) Sansone and Sansone (2010) for self-harm behaviour detection. Both are self-reported surveys where overall scores correspond to specific severity levels of the respective conditions.
@@
Increasingly, people turn to social media as a space to discuss their feelings and experiences and to find support Bucci et al. (2019); Naslund et al. (2016). Numerous initiatives have emerged to analyze social media content for health monitoring using NLP techniques, including CLPsych Tsakalidis et al. (2022) and eRisk Losada et al. (2017), through organized completion tasks. In this work, we use datasets from previous eRisk editions to validate our approach.
Recently, several studies have shown that closed-source LLMs struggle to achieve comparable results to state-of-the-art (SOTA) supervised methods in mental disorders classification tasks, both in zero-shot Amin et al. (2023) and few-shot XU et al. (2023) scenarios. We argue that it is possible to mitigate this issue by introducing an intermediate step where we instruct an LLM to answer questions from a standardized psychological questionnaire rather than directly requesting a diagnosis.
@@
To do that, we implement an adaptive RAG approach, combining retrieval and classification, to accurately predict users responses to psychological questionnaire items by analyzing their Reddit post history. Our goal is to test the capability of SOTA LLMs, both open- and closed-source, as potential annotators of standardized psychological questionnaires based on social media posts.
@@
The main contributions of this paper are: (1) We explore various combinations of open- and closed-source LLMs together with dense retrievers for predicting psychological questionnaires, evaluating how performance varies with different combinations of LLMs, prompt strategies, and retrieval models. (2) We compare our approach’s results with the best results obtained for the considered eRisk collections using primarily supervised models, showing how our unsupervised approach often outperforms the benchmarks. (3) We extend this paradigm to other mental disorders, introducing an interpretable and unsupervised method for predicting new behavioral disorders.
@@ 2 Related Works
Recent advancements in NLP have enabled the development of new and complex models across various areas, particularly in digital and mental health. In recent years, transformer-based models Vaswani et al. (2017) have emerged as powerful tools for predicting mental health disorders, leveraging vast amounts of textual data available from social media platforms. Pourkeyvan et al. (2024); Raj et al. (2024) employed different BERT models for depression detection on social media platforms, while Ji et al. (2022) developed MentalBERT, a specialized BERT model pre-trained on mental health-related content from Reddit, specifically designed to capture language patterns associated with mental disorders. Regarding LLMs, Yang et al. (2024a) introduced MentaLLaMA, an open source LLM specifically designed for interpretable mental health analysis on social media, building a multitask and multi-source instruction dataset to support model fine-tuning and evaluation. Furthermore, Yang et al. (2024b) presented a comprehensive analysis on how LLMs can enhance mental health care through real-time feedback and early intervention.
Concerning the use of NLP methods to predict psychological questionnaires’ responses, previous works Elourajini and Aïmeur (2022); Vu et al. (2020) used neural models to predict the 5 main personality traits Goldberg (2013) and the Myers-Briggs type indicator (MBTI) Schweiger (1985); Yang et al. (2021) based on user-generated textual posts and comments. Rosenman et al. (2024) propose an approach in which an LLM impersonates the interviewee, completing structured questionnaires, and then these responses are encoded as features to train a Random Forest model to predict the scores of the questionnaire’s items.
@@
With regard to eRisk data, recent approaches have been used to predict BDI-II responses using advanced computational methods. Pérez et al. (2023) implemented a computational approach capable of rapidly screening depressive disorders on social networks, using a retrieval pipeline and a combination of semi- and supervised models, implementing item-based classifiers to predict responses to different items. Recently, Ravenda et al. (2025) proposed a probabilistic neural architecture that uses a retrieval pipeline to select the most relevant posts for each user as input and predicts the final score using a Poisson distribution to better handle the ordinal nature of Likert scales. These latter two works differ from earlier approaches by incorporating a retrieval pipeline to filter the most relevant social media posts for each user and item before prediction, whereas previous methods fixed the number of posts per model, not considering their relevance. Additionally, while MBTI is a questionnaire with binary responses (binary classification), BDI-II uses a 4-point Likert scale, introducing an additional layer of complexity due to its ordinal nature.
Our work differs from previous approaches by focusing on a completely unsupervised scenario, leveraging LLMs in zero-shot contexts. The idea is to implement a retrieval pipeline to filter the most relevant posts and use LLMs to predict scores, linking the semantic content of posts to that of questionnaire items. To evaluate the effectiveness of LLMs in such tasks, we use two eRisk collections from the 2019 and 2020 editions Losada et al. (2019, 2020), which contain the post-history of Reddit users alongside their completed BDI-II questionnaires. After demonstrating the effectiveness of this unsupervised approach, we extend it to different mental disorders, highlighting the benefits of constraining LLM predictions to individual psychological questionnaire responses.
@@ 3 Research Questions
In the following, we state the main Research Questions of our work:
(RQ1.) Is it possible to fill a psychological questionnaire based on a user’s Reddit post history using a RAG-based approach in a completely unsupervised context? How does this approach compare in terms of performance to SOTA models for the considered datasets?
(RQ2.) How does the model effectiveness vary with changes in:
(RQ2a.) The LLM being used. To answer this question, we consider 4 LLMs: 2 open-source (Phi-3-mini and Phi-3.5-mini) and 2 closed-source (Claude-3.5-Sonnet and gpt-4o-mini).
@@
(RQ3.) Can our approach—completing a standardized psychological questionnaire to obtain a “psychological explanation” of why a user is associated with a certain risk levels—be extended to other mental disorders where no ground truth is available?
@@
Collection Disorder # of Users # of Posts Average # of Posts
Patient Control Patient Control Patient Control
eRisk 2017 Depression 52 349 18’706 217’665 359.7 623.7
eRisk 2019 Self-Harm 41 299 6’927 163’506 169.0 546.8
eRisk 2020 Self-Harm 104 319 11’691 91’136 112.4 285.6
Table 1: Users statistics of “early detection” task for eRisk 2017 (depression) and 2019, 2020 (self-harm) editions.
Collection Minimal Mild Moderate Severe # of Users # of Posts
eRisk 2019 4 4 4 8 20 10’380
eRisk 2020 10 23 18 19 70 33’600
Table 2: Users summary statistics of “Measuring the severity of depression” task for eRisk 2019 and 2020 editions.
Problem Definition. Computational approaches for predicting mental disorders from textual data, formulated as $f(\text{Text})\to Y$ Kim et al. (2020); Sekulić and Strube (2019), where Text represents textual information like social media posts or interview transcriptions, face limitations in interpretability and generalizability across contexts Paris et al. (2012); Friginal et al. (2017). These models, often neural, require large datasets to perform well. To address these limitations, we propose a method guided by psychological practice that leverages standardized questionnaires to assess the presence and/or intensity of mental disorders. The prediction is framed as a function correlating text and questionnaire items, $f(\text{Text},\text{Item}_{i})\to S_{i}$, where $S_{i}$ represents the user’s score for item $i$. The combination of these individual scores defines a final score used to diagnose symptoms as $\sum_{i}f(\text{Text},\text{Item}_{i})\to Y$. This approach not only provides an initial diagnosis but also enhances outcome’s interpretability by linking the final score to specific questionnaire items, offering a clear reasoning for the prediction. The proposed method comprises two main steps, illustrated in Figure 1: (1) retrieving the most relevant documents for each item using an adaptive dense retrieval approach; (2) generating responses in a zero-shot setting with LLMs, using the retrieved documents as context.
@@ 4.1 Datasets
For the analyses in this study, we use datasets from the eRisk collections of 2017, 2019, and 2020 Losada et al. (2017, 2019, 2020). The aim of this work is to evaluate the effectiveness of LLMs as assessors for mental disorders screening. For this purpose, we consider two different types of disorders: depression and self-harm, while focusing primarily on the former and using the latter as an additional proof-of-concept for our methodology. We use 5 datasets in total. For depression, we refer to task 1 of the 2017 edition and tasks 3 and 2 from the 2019 and 2020 editions. The first dataset includes social media posts and comments from users divided into two categories: those showing signs of depression and control users, belonging to the “early detection” task. The latter two datasets contain users post histories and responses to a standardized self-reported psychological questionnaire, the BDI-II, used for screening depressive disorders. These datasets belong to the “Measuring the severity of depression” task. Regarding self-harm, datasets contain social media posts and comments from users divided into two categories: users showing signs of self-harm and control users. Only posts preceding the user entry into the self-harm community are collected, in order to identify early signals and behavioral patterns that precede explicit help-seeking behavior, when intervention could be more effective and timely. The control group consists of both random users and users who actively participated in self-harm discussions (e.g., because they had a relative suffering from it). This dataset belongs to the “early detection” task. Table 1 shows some summary statistics for the datasets belonging to the early detection task, while Table 2 shows statistics for the Measuring the severity of depression task.
@@ 4.2 Adaptive Zero-Shot Retrieval Strategy
The number of Reddit posts per user can vary significantly, leading to two potential issues: exceeding the LLM token limit and reasoning degradation due to large input sizes Fraga (2024); Li et al. (2024). To manage this, we adopt a fully unsupervised posts retrieval approach using an embedding similarity approach, exploring 10 different dense retrieval models (see Table 3) and evaluating their effectiveness through LLM prediction accuracy.
@@
To account for variability in relevant documents per user, we employ the ABIDE-ZS method Ravenda et al. (2025). For each item, this approach identifies a neighborhood where the semantic meaning remains stable - i.e., the posts within this region share contextual relevance to the specific questionnaire’s item. Posts are retrieved based on semantic similarity, selecting the top $k^{*}$ posts, where $k^{*}$ is optimally determined by the ABIDE algorithm Noia et al. (2024) for each item. This allows us to avoid the need to fix a priori the number of $k$ documents to retrieve or specify a particular threshold (hence the name adaptive RAG for our approach).
@@
In Figure 1, we illustrate the process of retrieving Reddit posts for a specific user according to each questionnaire item. For a given user $i$, embeddings are created for each post (scatters) and each choice (denoted by the $\star$ symbol) within an item. Let $\mathbf{P}_{i}=\{\mathbf{p}_{i1},\ldots,\mathbf{p}_{im}\}$ represent the embeddings of the $m$ posts for user $i$, and $\mathbf{iq}_{j}=\{\mathbf{iq}_{j1},\mathbf{iq}_{j2},\mathbf{iq}_{j3},\mathbf{iq}_{j4}\}$ represent the embeddings of the 4 choices for item $j$. These item choices serve as queries (aka item-queries) to retrieve the most relevant posts from the user’s history. For each choice $\mathbf{iq}_{jl}$ (where $l=1,2,3,4$), the relevant Reddit posts $\mathbf{RRP}_{jl}$ (represented by colored dots in Figure 1) are retrieved by selecting the top $k^{*}$ posts with the highest embedding similarity scores: $\mathbf{RRP}_{ijl}=\{p_{i}\}_{p_{i}\in\aleph(iq_{jl})}$, where $\aleph(IC_{jl})$ represents the adaptive neighborhood of size $k^{*}$ for item $j$ and choice $l$. Specifically, for each item-query, we determine an adaptive neighborhood composed of $k^{*}$ Reddit posts nearest neighbors, where $k^{*}$ is not fixed but varies, from item to item, based on the local embeddings characteristics. Formally, we define two likelihood models: $M1$, which assumes different densities at the point $q$, representing a specific item-query, and its ($k+1$)-th neighbor, and $M2$, which assumes constant density. The log-likelihoods for these models are given by: $L_{M1}(q,k)=klog((k^{2})/(V_{q},k*V_{k}+1,k))-2k$ $L_{M2}(q,k)=2klog((2k)/(V_{q},k+V_{k}+1,k))-2k$ where $V_{q,k}$ is the volume of the hypersphere centered at $q$ with radius equal to the distance to its $k$-th neighbor. We then use a likelihood ratio test to determine the optimal $k^{*}(q)$: $k^{*}(q)=min(k:-2(L_{M2}(q,k)-L_{M1}(q,k))\geq D_{thr})$ where $D_{thr}$ is a threshold based on the desired confidence level. This adaptive approach allows us to identify a neighborhood size that is optimal for each query, adapting to local data densities and structures. In other words, the space identified by the adaptive neighborhood for each item-query can be viewed as a space where the semantics of the item-query remain constant, which is reflected in the region where the posts are semantically relevant to the item-query. Table 3 shows the statistics of the different dense retrieval models used, including the similarity measure adopted for each retriever (a more in depth discussion about retrievers used can be found in Appendix C). It can be observed that while the number of documents retrieved per item varies across models, no significant correlation exists between embedding dimension and the average number of retrieved documents (Pearson correlation, $p-value=0.70$). The impact of $k^{*}$ is discussed in Section D.1 of the Appendix.

### v2

@@ Abstract
In psychological practices, standardized questionnaires serve as essential tools for assessing mental health through structured, clinically-validated questions (i.e., items). While social media platforms offer rich data for mental health screening, computational approaches often bypass these established clinical assessment tools in favor of black-box classification. We propose a novel questionnaire-guided screening framework that bridges psychological practice and computational methods through adaptive Retrieval-Augmented Generation (aRAG). Our approach links unstructured social media content and standardized clinical assessments by retrieving relevant posts for each questionnaire item and using Large Language Models (LLMs) to complete validated psychological instruments. Our findings demonstrate two key advantages of questionnaire-guided screening: First, when completing the Beck Depression Inventory-II (BDI-II), our approach matches or outperforms state-of-the-art performance on Reddit-based benchmarks without requiring training data. Second, we show that guiding LLMs through standardized questionnaires can yield superior results compared to directly prompting them for depression screening, while also providing a more interpretable assessment by linking model outputs to clinically validated diagnostic criteria. Additionally, we show, as a proof-of-concept, how our questionnaire-based methodology can be extended to other mental conditions’ screening, highlighting the promising role of LLMs as psychological assessors. 111 Code available: https://github.com/Fede-stack/Adaptive-RAG-for-Psychological-Assessment
@@
Federico Ravenda1,2, Seyed Ali Bahrainian3,4, Andrea Raballo1, Antonietta Mira1,5, Noriko Kando2 federico.ravenda@usi.ch, bahrainian@brown.edu andrea.raballo@usi.ch, antonietta.mira@usi.ch, kando@nii.ac.jp 1Euler Institute, Università della Svizzera italiana, 2National Institute of Informatics 3Brown University, 4University of Tübingen, 5Insubria University
@@ 1 Introduction
According to the World Health Organization (WHO), one in seven adolescents experiences a mental health disorder Wiederhold (2022), with depression, anxiety, and behavioral disorders leading among young people. Following COVID-19, mental health conditions (MHCs) surged, with depressive disorders increasing by 28% in 2020 Kieling et al. (2011); Winkler et al. (2020). Given the extent of this need, the WHO Special Initiative for Mental Health prioritizes improving and expanding access to quality mental health interventions and services as a key strategic goal WHO (2022).
@@
In this work, we focus on different widely used standardized psychological questionnaires, specifically the Beck Depression Inventory-II (BDI-II) Beck (1996) for depression screening, the Self-Harm Inventory (SHI) Sansone and Sansone (2010) for self-harm behaviour detection, the SCOFF questionnaire Morgan et al. (1999) for eating disorders screening, and the Pathological Gambling Diagnostic Form from DSM-V Hopwood et al. (2012). All of them are self-reported surveys where overall scores correspond to specific severity levels of the respective conditions.
Despite their clinical validity, traditional psychological questionnaires present several practical limitations. They typically require in-person administration by trained professionals, making large-scale screening logistically challenging and resource-intensive. Access barriers disproportionately affect underserved populations, creating equity concerns in mental health assessment.
@@
Increasingly, people are turning to social networks as a space to express their feelings and experiences and find support Bucci et al. (2019); Naslund et al. (2016). Numerous initiatives have emerged to analyze social media content for health monitoring using NLP techniques, including CLPsych Tsakalidis et al. (2022) and eRisk Losada et al. (2017), through organized completion tasks.
Recent research has revealed significant limitations in using closed-source LLMs for mental disorder classification. Studies by Amin et al. (2023) and XU et al. (2023) demonstrate that both zero-shot and few-shot approaches struggle to match state-of-the-art (SOTA) supervised methods. These limitations stem from several challenges: (1) LLMs’ difficulty in directly mapping unstructured text to diagnostic categories, (2) the complex, multi-dimensional nature of mental health assessment requiring domain expertise, and (3) the semantic gap between social media language and clinical criteria. We propose mitigating these challenges through an intermediate step: rather than attempting direct diagnosis, we instruct LLMs to complete standardized psychological questionnaires, effectively decomposing the complex diagnostic task into structured, clinically-validated assessment items.
@@
To achieve that, we propose an adaptive RAG approach (aRAG), combining retrieval and classification, to accurately predict users’ responses to psychological questionnaire items by analyzing their Reddit post history. Unlike existing methods that use fixed retrieval parameters or direct classification, our approach automatically determines the optimal number of relevant posts needed for each questionnaire item, adapting to the semantic density and relevance of available content. To the best of our knowledge, we demonstrate, for the first time, how LLMs (both open- and closed-source) can serve as effective annotators of standardized psychological questionnaires by analyzing social media posts through aRAG.
@@
The main contributions of this paper are: (1) We explore various combinations of open- and closed-source LLMs together with dense retrievers for predicting psychological questionnaires, evaluating how performance varies with different combinations of LLMs, prompt strategies, and retrieval models. (2) We compare the results of our approach with the best results obtained for the considered eRisk collections using primarily supervised models, showing how our unsupervised approach often outperforms the benchmarks. (3) We extend this paradigm to other mental disorders, introducing an interpretable and unsupervised method for predicting new MHCs.
These results confirm that LLMs can serve as effective and promising psychological assessors when their predictions are guided by standardized clinical instruments, bridging the gap between language models and psychometric practice.
@@ 2 Related Works
Recent advancements in NLP have enabled the development of new and complex models across various areas, particularly in digital and mental health Ríssola et al. (2019). Transformer-based models Vaswani et al. (2017) have significantly advanced mental health analysis on social media platforms. While initial work used BERT variants for depression detection Raj et al. (2024); Ríssola et al. (2020), specialized models like MentalBERT Ji et al. (2022) emerged, pre-trained specifically on mental health-related content. A novel direction explored how emotion manifests in depressed individuals’ Bucur et al. (2022) and a broader spectrum of mental disorders De Grandi et al. (2024) from social media posts, using emotional and psychological markers to provide interpretable assessment.
Recently, Large Language Models have shown increasing promise, with approaches like MentaLLaMA Yang et al. (2024a) offering interpretable analysis, and Varadarajan et al. (2024) combining theoretical frameworks with computational techniques for suicide risk assessment. These advances suggest potential for real-time intervention and support Yang et al. (2024b).
Concerning the use of NLP methods to predict psychological questionnaires’ responses, early approaches used neural models to predict personality traits Elourajini and Aïmeur (2022) and Myers-Briggs indicators Yang et al. (2021) from user-generated social media content. BERT embeddings have been leveraged to link social media expressions with psychological assessments Vu et al. (2020); Atari et al. (2023). More recently, Rosenman et al. (2024) used an LLM to impersonate interviewees and complete questionnaires, using these responses as features in a Random Forest to predict new questionnaire scores.
@@
With regard to eRisk data, recent approaches were used to predict BDI-II responses using advanced computational methods. Pérez et al. (2023) introduced a retrieval-based framework with item-based classifiers for depression screening, while Ravenda et al. (2025) proposed a probabilistic approach combining retrieval mechanisms to handle ordinal Likert scales. Both approaches innovate via retrieval-based selection of relevant social media content, departing from previous methods that used fixed post selection.
Our work differs from previous approaches by focusing on a completely unsupervised scenario, leveraging LLMs through an aRAG approach, filtering the most relevant posts and use LLMs to predict scores, establishing a semantic mappings between social media content and standardized questionnaire items. To evaluate the effectiveness of our approach in such tasks, we use two eRisk collections from the 2019 and 2020 editions Losada et al. (2019, 2020), which contain the post-history of Reddit users alongside their completed BDI-II questionnaires. After demonstrating the effectiveness of this unsupervised approach, we extend it to different MHCs, highlighting the benefits of constraining LLM predictions to individual psychological questionnaire responses.
@@ 3 Research Questions
The main Research Questions we address are:
(RQ1.) Is it possible to fill out a psychological questionnaire automatically based on a user’s Reddit post history in an unsupervised context? How does this compare in terms of performance to SOTA models for the considered datasets?
(RQ2.) How does the model’s effectiveness vary with changes in:
(RQ2a.) The LLM being used. To answer this question, we consider six LLMs: two large-scale open-source (Qwen 2.5 70B, DeepSeek V3), two lightweight open-source (Phi-3-mini, Phi-3.5-mini), and 2 closed-source (Claude-3.5-Sonnet, gpt-4o-mini).
@@
(RQ3.) Can our approach -completing a standardized psychological questionnaire to obtain a “psychological explanation” of why a user is associated with certain risk levels - be extended to other MHCs without training data?
@@
Problem Definition. While existing computational approaches typically frame mental health prediction as a direct mapping from text to diagnosis $f(\text{Text})\to Y$ Kim et al. (2020); Sekulić and Strube (2019), where Text represents textual information like social media posts or interview transcriptions, this black-box formulation faces significant challenges in clinical applicability, interpretability, and generalizability across contexts Paris et al. (2012); Friginal et al. (2017). These models, often neural, require large datasets to perform well. Our work introduces a novel paradigm that bridges computational and clinical practice by leveraging standardized psychological questionnaires as an intermediate structured representation. By reformulating the prediction task through questionnaire items, our method offers several key advantages: (1) clinical interpretability through standardized assessment criteria, (2) transparent reasoning through item-level predictions, and (3) alignment with established psychological practice. The prediction is framed as a function linking text and questionnaire items, $f(\text{Text},\text{Item}_{i})\to S_{i}$, where $S_{i}$ represents the user’s score for item $i$. The combination of these individual scores defines a final score used to diagnose symptoms as $\sum_{i}f(\text{Text},\text{Item}_{i})\to Y$. The proposed method comprises two main steps, illustrated in Figure 1: (1) retrieving the most relevant posts for each item using an adaptive dense retrieval approach; (2) generating responses in a zero-shot setting with LLMs, using the retrieved documents as context.
@@ 4.1 Datasets
Our analyses uses seven datasets from the 2017, 2019, 2020, and 2022 eRisk collections Losada et al. (2017, 2019, 2020); Parapar et al. (2022), encompassing different MHCs. The depression datasets consist of two distinct tasks. The “early detection task dataset” from eRisk 2017 (used in Section 5.3) provides binary classification data (depression vs. control users) through user posts and comments, with users labeled based on signs of depression. The “severity assessment task datasets” from eRisk 2019 and 2020 (used in Section 5.1) contain user post histories along with their responses to the BDI-II questionnaire, enabling detailed depression severity measurement. The self-harm, anorexia, and pathological gambling datasets belong to the “early detection tasks datasets” from eRisk 2019, 2020, and 2022 (used in Section 5.2). These datasets contain posts preceding users’ entry into self-harm, anorexia, and gambling communities, aiming to identify early warning signals before explicit help-seeking behavior. Table 1 summarizes the user and post distributions across conditions for the early detection task datasets, while Table 2 presents detailed statistics for the severity assessment task datasets, including the distribution of depression severity levels.
@@ 4.2 Adaptive Zero-Shot Retrieval Strategy
In this subsection, we address the challenge of retrieving relevant social media content for psychological assessment. The number of Reddit posts per user varies widely, leading to potential issues such as exceeding the LLM token limit and reasoning degradation due to large inputs Fraga (2024); Li et al. (2024, 2025). To mitigate these issues, we adopt a fully unsupervised retrieval strategy based on embedding similarity, exploring 10 different dense retrieval models (see Table 3) and evaluating their effectiveness through LLM prediction accuracy.
Collection # of Users # of Posts
Patient Control Patient Control
eRisk 2017 - Depression 52 349 359.7 623.7
eRisk 2019 - Self-Harm 41 299 168.9 212.4
eRisk 2020 - Self-Harm 104 319 112.4 285.7
eRisk 2019 - Anorexia 73 742 241.4 745.1
eRisk 2022 - Gambling 81 1998 180.6 507.6
Table 1: Users statistics of “early detection task datasets” for eRisk 2017, 2019, 2020, and 2022 for depression, self-harm, anorexia, and pathological gambling.
Collection Minimal Mild Moderate Severe
eRisk 2019 4 4 4 8
# of Users: 20
# of Posts: 10’380
eRisk 2020 10 23 18 19
# of Users: 70
# of Posts: 33’600
Table 2: Users summary statistics of “severity assessment task datasets” for eRisk 2019 and 2020 editions.
@@
To account for variability in relevant posts per user, we employ the ABIDE-ZS method Ravenda et al. (2025). For each item, this approach identifies a neighborhood where the semantic meaning remains stable - i.e., the posts within this region share contextual relevance to the specific questionnaire’s item. Posts are retrieved based on semantic similarity, selecting the top $k^{*}$ posts, where $k^{*}$ is optimally determined by the ABIDE algorithm Noia et al. (2024) for each item (see Section B). This eliminates the need to fix a priori the number of $k$ posts to retrieve or set a threshold, making our approach adaptive.
@@
Figure 1 illustrates the retrieval process. For a each user $i$, embeddings are generated for their posts (scatters) and for the four response choices (denoted by the $\star$ symbol) within an item. Let $\mathbf{P}_{i}=\{\mathbf{p}_{i1},\ldots,\mathbf{p}_{im}\}$ represent the embeddings of the $m$ posts for user $i$, and $\mathbf{iq}_{j}=\{\mathbf{iq}_{j1},\mathbf{iq}_{j2},\mathbf{iq}_{j3},\mathbf{iq}_{j4}\}$ represent the embeddings of the four choices for item $j$. These item choices serve as queries (aka item-queries) to retrieve the most relevant posts from the user’s history. For each choice $\mathbf{iq}_{jl}$ (where $l=1,2,3,4$), the relevant Reddit posts $\mathbf{RRP}_{jl}$ (represented by colored dots in Figure 1) are retrieved by selecting the top $k^{*}$ posts with the highest embedding similarity scores: $\mathbf{RRP}_{ijl}=\{p_{i}\}_{p_{i}\in\aleph(iq_{jl})}$, where $\aleph(IC_{jl})$ represents the adaptive neighborhood of size $k^{*}$ for item $j$ and choice $l$. Consider user posts and questionnaire item choices embedded in ${R}^{D}$, where $D$ is the dimension of the embeddings. The optimal $k^{*}$ is determined by the local density of posts around each query point, ensuring we retrieve exactly the number of posts needed to maintain semantic coherence. In other words, the space identified by the adaptive neighborhood for each item-query can be seen as a space where the semantics of the item-query remain constant, which is reflected in the region where the posts are semantically relevant to the item-query. Table 3 shows the statistics of the different dense retrieval models used, including the similarity measure adopted for each retriever (a more in-depth discussion about retrievers used can be found in Appendix D). We observe that while the number of documents retrieved per item varies across models, no significant correlation exists between embedding dimension and the average number of retrieved documents (Pearson correlation, $p-value=0.70$). The impact of $k^{*}$ is discussed in Section E.2 of the Appendix.